Block B has a mass of 3.80kg and is moving to the left at a speed of 2.20m/s. Block A has a mass of 3.00kg and is moving to the right. The two blocks undergo a perfectly inelastic collision. What should the velocity of the Block A be in order to have the two blocks move to the left with a speed of 0.37 m/s after the collision?
Answer #1

Let us consider the left side as positive and the right side as negative.

Mass of block A = m1 = 3 kg

Mass of block B = m2 = 3.8 kg

Speed of block A before the collision = V1

Speed of block B before the collision = V2 = 2.2 m/s

It is a perfectly inelastic collision that is after the collision both the blocks will stick together which means they will move at the same speed after the collision.

Speed of the blocks after the collision = V3 = 0.37 m/s

By conservation of momentum,

m1V1 + m2V2 = (m1 + m2)V3

(3)V1 + (3.8)(2.2) = (3 + 3.8)(0.37)

V1 = -1.948 m/s

The negative sign indicates that block A is moving to the right.

Velocity of block A before the collision = 1.948 m/s to the right